$SPAD/src/algebra product.spad

The Axiom Team

July 17, 2011

Abstract

This domain implements cartesian product for a pair of (possibly different) domains. If the underlying domains are both Finite then the resulting Product is also Finite and can be enumerated via size(), index(), location(), etc. The index of the second component (B) varies most quickly.

Contents

1 domain PRODUCT Product 3 2 License 5

1 domain PRODUCT Product

⟨domain PRODUCT Product⟩≡

)abbrev domain PRODUCT Product
++ Description:
++ This domain implements cartesian product
Product (A:SetCategory,B:SetCategory) : C == T
where
C == SetCategory  with
if A has Finite and B has Finite then Finite
if A has Monoid and B has Monoid then Monoid
if A has AbelianMonoid and B has AbelianMonoid then AbelianMonoid
if A has CancellationAbelianMonoid and
   B has CancellationAbelianMonoid then CancellationAbelianMonoid
if A has Group  and B has Group  then  Group
if A has AbelianGroup and B has AbelianGroup then  AbelianGroup
if A has OrderedAbelianMonoidSup and B has OrderedAbelianMonoidSup
                                      then OrderedAbelianMonoidSup
if A has OrderedSet and B has OrderedSet then  OrderedSet
makeprod     : (A,B) -> %
 ++ makeprod(a,b) \undocumented
selectfirst  :   %   -> A
 ++ selectfirst(x) \undocumented
selectsecond :   %   -> B
 ++ selectsecond(x) \undocumented
T == add
--representations
Rep := Record(acomp:A,bcomp:B)
--declarations
x,y: %
i: NonNegativeInteger
p: NonNegativeInteger
a: A
b: B
d: Integer
--define
coerce(x):OutputForm == paren [(x.acomp)::OutputForm,
                               (x.bcomp)::OutputForm]
x=y ==
    x.acomp = y.acomp => x.bcomp = y.bcomp
    false

makeprod(a:A,b:B) :%   == [a,b]
selectfirst(x:%) : A   == x.acomp
selectsecond (x:%) : B == x.bcomp
if A has Monoid and B has Monoid then
1 == [1$A,1$B]
x * y == [x.acomp * y.acomp,x.bcomp * y.bcomp]
x ** p == [x.acomp ** p ,x.bcomp ** p]
if A has Finite and B has Finite then
size == size$A * size$B
index(n) == [index((((n::Integer-1) quo size$B )+1)::PositiveInteger)$A,
             index((((n::Integer-1) rem size$B )+1)::PositiveInteger)$B]
random() == [random()$A,random()$B]
lookup(x) == ((lookup(x.acomp)$A::Integer-1) * size$B::Integer + lookup(x.bcomp)$B::Integer)::PositiveInteger
hash(x) == hash(x.acomp)$A * size$B::SingleInteger + hash(x.bcomp)$B
if A has Group and B has Group then
inv(x) == [inv(x.acomp),inv(x.bcomp)]
if A has AbelianMonoid and B has AbelianMonoid then
0 == [0$A,0$B]
x + y == [x.acomp + y.acomp,x.bcomp + y.bcomp]
c:NonNegativeInteger * x == [c * x.acomp,c*x.bcomp]
if A has CancellationAbelianMonoid and
B has CancellationAbelianMonoid then
  subtractIfCan(x, y) : Union(%,"failed") ==
    (na:= subtractIfCan(x.acomp, y.acomp)) case "failed" => "failed"
    (nb:= subtractIfCan(x.bcomp, y.bcomp)) case "failed" => "failed"
    [na::A,nb::B]
if A has AbelianGroup and B has AbelianGroup then
- x == [- x.acomp,-x.bcomp]
(x - y):% == [x.acomp - y.acomp,x.bcomp - y.bcomp]
d * x == [d * x.acomp,d * x.bcomp]
if A has OrderedAbelianMonoidSup and B has OrderedAbelianMonoidSup then
sup(x,y) == [sup(x.acomp,y.acomp),sup(x.bcomp,y.bcomp)]
if A has OrderedSet and B has OrderedSet then
x < y ==

               xa:= x.acomp ; ya:= y.acomp
               xa < ya => true
               xb:= x.bcomp ; yb:= y.bcomp
               xa = ya => (xb < yb)
               false
--     coerce(x:%):Symbol ==
--      PrintableForm()
--      formList([x.acomp::Expression,x.bcomp::Expression])$PrintableForm

2 License

⟨license⟩≡

--Copyright (c) 1991-2002, The Numerical ALgorithms Group Ltd.
--All rights reserved.
--
--Redistribution and use in source and binary forms, with or without
--modification, are permitted provided that the following conditions are
--met:
--
--    - Redistributions of source code must retain the above copyright
--      notice, this list of conditions and the following disclaimer.
--
--    - Redistributions in binary form must reproduce the above copyright
--      notice, this list of conditions and the following disclaimer in
--      the documentation and/or other materials provided with the
--      distribution.
--
--    - Neither the name of The Numerical ALgorithms Group Ltd. nor the
--      names of its contributors may be used to endorse or promote products
--      derived from this software without specific prior written permission.
--
--THIS SOFTWARE IS PROVIDED BY THE COPYRIGHT HOLDERS AND CONTRIBUTORS "AS
--IS" AND ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, BUT NOT LIMITED
--TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A
--PARTICULAR PURPOSE ARE DISCLAIMED. IN NO EVENT SHALL THE COPYRIGHT OWNER
--OR CONTRIBUTORS BE LIABLE FOR ANY DIRECT, INDIRECT, INCIDENTAL, SPECIAL,
--EXEMPLARY, OR CONSEQUENTIAL DAMAGES (INCLUDING, BUT NOT LIMITED TO,
--PROCUREMENT OF SUBSTITUTE GOODS OR SERVICES; LOSS OF USE, DATA, OR
--PROFITS; OR BUSINESS INTERRUPTION) HOWEVER CAUSED AND ON ANY THEORY OF
--LIABILITY, WHETHER IN CONTRACT, STRICT LIABILITY, OR TORT (INCLUDING
--NEGLIGENCE OR OTHERWISE) ARISING IN ANY WAY OUT OF THE USE OF THIS
--SOFTWARE, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGE.

⟨* ⟩≡ ⟨license⟩

⟨domain PRODUCT Product⟩

References^{[1] nothing}

7